### v1

# Disorder to Order Transition in 1D Nonreciprocal Cahn-Hilliard Model

Abstract
We extensively study the phenomenology of one dimensional Nonreciprocal Cahn Hilliard model for varying nonreciprocity $(\alpha)$ and different boundary conditions. At small $\alpha$, a perturbed uniform state evolves to defect laden configuration that lack global polar order. Defects are the sources and sinks of travelling waves and nonreciprocity selects defects with a unique wave number that increases monotonically with $\alpha$. A critical threshold $\alpha_{c}$ marks the onset of a transition to states with finite global polar order. For periodic boundaries, above $\alpha_{c}$, the system shows travelling waves that are completely ordered. In contrast, travelling waves are incompatible with Dirichlet and Neumann boundaries. Instead, for $\alpha\gtrsim\alpha_{c}$, we find fluctuating domains that show intermittent polar order and at large $\alpha$, the system partitions into two domains with opposite polar order.
I Introduction
The Nonreciprocal Cahn-Hilliard (NRCH) model describes phase separation of multi-component mixtures in the presence of nonreciprocal couplings [1, 2, 3, 4, 5, 6, 7, 8]. Systems with nonreciprocity are intrinsically out of equilibrium and they show rich phenomenology [9, 10, 11, 12, 13, 14, 15, 16]. In the NRCH model, the spontaneous breaking of parity and time-reversal symmetries leads to the formation of travelling density bands [1, 2], suppressed coarsening dynamics [17, 1], localised states [18], chaotic steady states [8], true long-range polar order in two dimensions [7], and enhanced stability in multispecies mixtures [5]. Introduced phenomenologically, the NRCH model emerges as a universal amplitude equation for a conserved-Hopf instability in systems with two conservation laws [19] and can also be derived from a systematic coarse-graining of a microscopic model of phoretically active Janus colloids [20, 21].
In Rana and Golestanian [3, 4], we presented defect dynamics of the NRCH model in two dimensions (2D). The model admits spirals with unit magnitude topological charge and topologically neutral targets. These defects are the sources of the travelling waves, and thus the precursors of global polar order. For a given strength of nonreciprocal coupling $\alpha$, defects with a unique asymptotic wave number, $k_{\infty}\propto\sqrt{\alpha}$ and amplitude $R_{\infty}=\sqrt{1-k_{\infty}^{2}}$, are selected [3]. At a critical threshold of nonreciprocity $\alpha_{c}$; we find a disorder to order transition. For $\alpha<\alpha_{c}$, random disordered states evolve to quasi-stationary defect networks with no global polar order. On the other hand, for $\alpha>\alpha_{c}$, we find travelling waves. Below $\alpha_{c}$, the topological composition of the defect networks is also $\alpha$ dependent. For $\alpha\ll\alpha_{c}$, we always find spirals, whereas for $\alpha\lesssim\alpha_{c}$ we primarily observe target networks. Our further study [4] on the interactions between a pair of defects revealed that the stability of targets fundamentally alters the nature of defect interactions and sets them apart from the previously studied out of equilibrium systems with a nonconserved order parameter [22, 23, 24, 25]. Thus, NRCH provides a fertile ground to investigate the defect phenomenology in conserved nonreciprocal systems.
In this paper, we present an extensive study on the defects of the NRCH model in one dimension (1D). Our main motivation is to investigate defect dynamics of the NRCH model in a simpler setting as compared to our previous 2D studies [3, 4]. Charged spirals are ruled out, and we observe stable sources and sinks, which are the direct analogue of two dimensional targets. These sources and sinks are arranged in an alternating configurations and their motion is limited to the line. Similar to 2D, sources and sinks with a particular wavenumber $k_{\infty}$ that increases with $\alpha$ are selected. The 1D NRCH model also shows a disorder-order transition. Crucially and in a stark contrast to 2D, this transition occurs at the crossover $\alpha$ predicted by the Eckhaus instability of the plane waves. We further investigate the features of the disorder-order transition for different boundary conditions. Since travelling waves are not accepted solutions of the Neumann or the Dirichlet boundary conditions, at $\alpha\gtrsim\alpha_{\times}$, the system shows intermittent behaviour, with fluctuating polar order for these boundary conditions. For larger $\alpha$, the system splits into subdomains with opposing polar order.
The rest of the paper is organized as follows. In Section II, we present our model and methods. In Section III, we present our results for periodic boundary conditions. We highlight the features of the source and sink solutions, wave number selection, defect density, polar order, and the disorder to order transition. In Section IV, we consider Dirichlet and Neumann boundary conditions and discuss how the boundary conditions alter the phenomenology. We conclude the paper with a discussion and future perspectives.
II Model and Methods
We consider a minimal model of two conserved scalar fields $\phi_{1}(x,t)$ and $\phi_{2}(x,t)$ with nonreciprocal interactions on a one dimensional domain. The conserved dynamics of the complex scalar order parameter $\phi=\phi_{1}+i\phi_{2}$ is governed by the continuity equation $\partial_{t}\phi=\partial_{x}^{2}\mu$, where $\mu=\delta\mathcal{F}/\delta\phi^{*}+i\alpha\phi$ is the non-equilibrium chemical potential. The equilibrium contribution to $\mu$ promotes phase-separation and is derived from the free energy functional $\mathcal{F}=\int\mathop{}\!\mathrm{d}x\left(-|\phi|^{2}/2+|\phi|^{4}/4+|\nabla\phi|^{2}/2\right)$. The nonequilibrium part encodes the nonreciprocal interactions between the two species; quantified by the parameter $\alpha$. For $\alpha=0$, the model describes equilibrium phase-separation of two interacting species, and an initial disordered state coarsens towards a bulk-phase separated state. Any $\alpha\neq 0$ sets the system out of equilibrium and it describes nonreciprocal phase-separation. A positive $\alpha$ implies that $\phi_{1}$ chases $\phi_{2}$, and the model is invariant under the simultaneous transformation $\alpha\to-\alpha$ and $\phi_{1,2}\to\phi_{2,1}$. Combining everything together, we obtain the following non-dimensional equation for the evolution of $\phi(x,t)$ [3, 4, 1, 2]
$\partial_{t}{\phi}$ $=\partial_{x}^{2}\left[(-1+i\alpha)\phi+|\phi|^{2}\phi-\partial_{x}^{2}\phi\right].$ (1)
The homogeneous state $\phi=0$ is unstable to small perturbations and a linear stability yields that perturbations of the form $\delta\phi(k,t)\exp(ikx)$ evolve as
$\partial_{t}{\delta\phi}(k,t)=k^{2}(1-i\alpha-k^{2})\delta\phi(k,t),$ (2)
which reveals that $\forall~|k|<1$ the perturbations grow in an oscillatory manner. The behaviour of the system depends on $\alpha$, domain length $L$, and the choice of boundary conditions. In this manuscript, we systematically study the phase-behaviour of Eq. 1 with varying strength of nonreciprocity $(\alpha)$ for various domain sizes $(L)$ and boundary conditions. We consider three different kinds of boundary conditions that are relevant to the phenomenon of phase separation [26]:
  1. (i)
    Periodic boundary conditions (P-BC), where
    $\phi(x+L)=\phi(x)~\mathrm{and}~\mu(x+L)=\mu(x).$
  2. (ii)
    No-flux boundary conditions (N-BC), where
    $\partial_{x}\phi|_{0}=\partial_{x}\phi|_{L}=0~\mathrm{and}~\partial_{x}\mu|_{0}=\partial_{x}\mu|_{L}=0.$
  3. (iii)
    Dirichlet boundary conditions (D-BC), where
    $\phi(0)=\phi(L)=0~\mathrm{and}~\mu(0)=\mu(L)=0.$
To study the phenomenology, we numerically integrate Eq. 1 on a 1D domain of length $L$ that is discretized using $N$ points. For P-BC, we use a pseudo-spectral algorithm that employs a second-order exponential time-differencing scheme [27] for time marching. For N-BC and D-BC, we perform the simulations using a Chebyshev polynomials based spectral Tau method implemented using the Dedalus framework [28]. Additionally, we also use Dedalus to cross-check our results for the periodic domain. The source code to simulate the 1D NRCH model using Dedalus is available at https://github.com/navdeeprana/DedalusScripts.
We start all our simulations with a small perturbation to the homogeneous state. i.e., $\phi(x,0)=\delta\phi(x)$, where at every point $x$ the initial perturbation $\delta\phi(x)$ is drawn from a uniform distribution. Further, we impose $\int_{0}^{L}\phi(x,0)\mathop{}\!\mathrm{d}x=0$, which remains constant throughout the entire evolution due to the overall conservation law. Unless reported otherwise, we average over eight or more independent realizations of initial conditions for each data point reported in the paper.
Figure 1: Kymograph plot of the polar order parameter $J(x,t)$ highlighting the evolution of disordered states. Defects are located on zeroes of $J(x,t)$, which appear white in the kymograph. We also mark the exact defect positions (small black points) at late times. Initially, numerous source-sink pairs are formed, which merge and the system settles into a stable configuration, which persists till the end of the simulations.
Figure 2: Sources and sinks in the 1D NRCH model. (a, b) Representative plot of $\real{\phi}$ and the polar order parameter $J$ (normalised by it’s maximum value), respectively. Sources (blue circles) and sinks (orange circles) are arranged in an alternating manner on the 1D domain. $J$ vanishes at defect centres and is constant far from defects, where the wavefront is that of a plain wave. (c, d) Plot of $R(x)$ and $k(x)$ for sources (left column) and sinks (right column) for different values of $\alpha$. For these plots, $x$ is measured from the centre of the defects. At defect centres, $\partial_{x}R$ and $k(x)$ vanish. Far from the defects, $R(x)$ and $k(x)$ approach their plain wave forms.
III Periodic boundary conditions
For P-BC, the NRCH model Eq. 1 admits travelling wave solutions of the form
$\phi(x,t)=R\,e^{i(kx-\omega t)},$ (3)
where $k<1$, $R^{2}=1-k^{2}$, and $\omega=\alpha k^{2}$. For small $k$, travelling states are stable to small perturbations, however an Eckhaus instability kicks in at large $k$ and restricts the allowed wave number range to $0\leq k^{2}<1/3$ [23, 8, 3, 29]. In our previous studies [3, 4], we have shown that in 2D, at low $\alpha$, disordered states evolves to defect configurations composed of spirals and targets. In 1D, we find similar phenomenology. A random state evolves into multi-defect configurations that persist up to the end of simulation time (see Fig. 1 and Fig. 2). In what follows, we discuss the properties of these states in detail.
III.1 Sources and Sinks
The 1D NRCH model Eq. 1 admits solutions of the form
$\phi(x,t)=R(x)\,e^{i\left[Z(x)-\omega t\right]},$ (4)
where $x$ is measured from defect location, $R(x)$ is the amplitude, $Z(x)$ is the phase, and $k(x)\equiv\mathop{}\!\mathrm{d}Z/\mathop{}\!\mathrm{d}x$. They are the sources and sinks of travelling waves and we shall collectively call them defects. In Fig. 2(a), we plot a representative multi-defect configuration obtained from our P-BC simulations. The sources and sinks are arranged in an alternating manner on the 1D domain. Furthermore, for a periodic domain, the total number of sources is always equal to the total number of sinks. It is easy to identify the position of these defects by using the polar order parameter
$J(x,t)=\mathrm{Im}(\phi^{*}\partial_{x}\phi)=\frac{1}{2i}\left(\phi^{*}\partial_{x}\phi-\phi\partial_{x}\phi^{*}\right).$ (5)
As shown in Fig. 2(b), at a source (sink), $J$ vanishes with a positive (negative) slope, thus at a given time $t$, the number of defects is equal to the number of zeroes of $J(x,t)$. Substituting the defect ansatz in the definition of $J(x,t)$, we further find that at the defect core $\partial_{x}(R^{2})=\partial_{x}(\phi\bar{\phi})=2\bar{\phi}\partial_{x}\phi=2\phi\partial_{x}\bar{\phi}$. Thus, at the defect positions,
$\partial_{x}R(x)=0,~\mathrm{and}~k(x)=0.$ (6)
In Fig. 2(c,d), we plot the profiles of $R(x)$ and $k(x)$ for sources and sinks for different $\alpha$, where $x$ is measured from the location of the defect. For a source, $k(x)>0(<0)$ when $x>0(<0)$, and vice versa for a sink. It is evident that R(x)=R(x)R(-x)=R(x) and k(x)=k(x)k(-x)=-k(x), which implies that for small $x$, $R(x)\sim a_{0}-a_{2}x^{2}$ and $k(x)\sim b_{1}x-b_{3}x^{3}$. Far away from the defect centre, the wavefront approaches that of a travelling wave (see Fig. 2(a)). Thus, when $x\to\pm\infty$, $k(x)\to\pm k_{\infty}$ for a source and $k(x)\to\mp k_{\infty}$ for a sink. The amplitude $R(x)\to\sqrt{1-k_{\infty}^{2}}$ also becomes constant for both and to ensure proper oscillations, we require $\omega=\alpha k_{\infty}^{2}$. Note that the sources are the direct 1D analogue of targets [3, 4] and topologically charged spirals do not exist in 1D.
Figure 3: Wave number selection in 1D. For a given $\alpha$, sources and sinks with a particular wave number are selected. Using a least-square fit to power law forms, we obtain $k_{\infty}\sim\alpha^{0.6}$ (Solid black line). The selected wave number is comparable to the $k_{\infty}$ obtained for the targets in 2D (Orange circles) [3]. Inset: $k_{\infty}$ for a single source-sink pair is identical to the $k_{\infty}$ for a multi-defect disordered configuration.
Figure 4: (a) Defect density $\rho_{D}$ versus $\alpha$ for P-BC shows nontrivial behaviour. It is low at certain values of $\alpha\simeq 0.08,0.2,0.3$, which we call “resonances” (shaded green). Overall it first increases with $\alpha$, peaks around $\alpha\sim 0.2$ and then starts to decrease. Above $\alpha_{c}\sim 0.6$, $\rho_{D}$ vanishes marking the onset of the travelling wave phase. $\alpha_{c}$ is close to the crossover threshold $\alpha_{\times}\sim 0.62$ predicted by the wavenumber selection. (b) Average polar order $\bar{J}$ versus $\alpha$ in the long time steady state. For small $\alpha$, $\bar{J}\sim 0$, at large $\alpha$, $\bar{J}\sim 1$. Near $\alpha_{c}$ (grey shaded region), $\bar{J}$ shows large fluctuations due to finite size effects. Inset: The fraction of simulations $(F_{T})$ that show defect states for $\alpha$ close to $\alpha_{c}$. Below $\alpha_{c}$, $F_{T}=1$, above $\alpha_{c}$, all simulations transition to travelling waves, thus $F_{T}=0$. (c) Time ($t_{\text{st}}$) after which the system reaches the steady-state value of $\rho_{D}$. At resonances, defect merger events continue for long times, thus $t_{\text{st}}$ is orders of magnitude larger. To verify that the resonances are not the effect of domain size, we also plot results for $L=256\pi$ (black points), which shows similar behaviour. (d) Minimum inter-defect separation $d_{\mathrm{min}}$ versus $\alpha$. As defects keep merging for long times, we find sharp jumps in $d_{\mathrm{min}}$ in the resonance regions, which we also observe in 2D [4]. All plots share the same legend keys, and the location of the three resonances that are shaded green are at the same values of $\alpha\simeq 0.08,0.2,0.3$ for panels (a), (c), and (d).
III.2 Wave number selection
A we show in Fig. 3, for a given $\alpha$, sources and sinks with a particular value of the asymptotic wave number $k_{\infty}$ are selected. $k_{\infty}$ increases monotonically with $\alpha$, and from a least-square fit, we obtain $k_{\infty}\sim\alpha^{0.6}$. Further, $k_{\infty}$ is comparable to the wave number selected for targets in 2D. In Fig. 3(inset), we verify that $k_{\infty}$ for a single source-sink pair configuration is the same as the $k_{\infty}$ obtained from multi-defect configurations that evolved from disordered initial conditions. A monotonically increasing $k_{\infty}$ predicts a crossover value of non reciprocity, $\alpha_{\times}$, such that when $\alpha\to\alpha_{\times}$, $k_{\infty}\to 1$, and $R_{\infty}\to 0$. It implies that the defect solutions will cease to exist for $\alpha\geq\alpha_{\times}$. Using a least-square fit, we find $\alpha_{\times}\sim 1.5$. However this prediction of $\alpha_{\times}$ is not entirely correct as it does not account for the Eckhaus instability. Since the plane waves generated by the defects are unstable for $k^{2}\geq 1/3$ [23, 8, 3, 29], we expect that the defect solutions will vanish for $k_{\infty}^{2}\geq 1/3$ which yields $\alpha_{\times}\sim 0.62$. Thus wavenumber selection and the Eckhaus instability restrict the defect solutions to the range $0<\alpha<0.62$.
III.3 Defect density and the Global polar order
We now focus on the statistical properties of the multi-defect configurations. Using the fact that $J=0$ at a source or a sink, we define the 1D defect density $\rho_{D}$ simply as the number of zeros of $J$ per unit length. In Fig. 4(a), we plot $\rho_{D}$ in the steady state for different values of $\alpha$. Since $\alpha=0$ is the equilibrium limit of the model, $\rho_{D}$ vanishes as $\alpha\to 0$. Furthermore, we find that defects cease to exist above $\alpha_{c}\sim 0.6$ and consequently $\rho_{D}$ vanishes, instead we find travelling waves. This marks the onset of disorder-to-order transition that was also observed in 2D [3]. As shown in Fig. 4(b), we plot the average polar order $\bar{J}$ for different values of $\alpha$, in the steady state. $\bar{J}$ is defined as
[…]
where $\left$$\ldots\right$> implies averaging over space and time in the steady-state, and we use the dominant wave number of the patterns $q_{p}$ for normalization. For travelling waves, $J=R^{2}k=k-k^{3}$, and $q_{p}=k$, which sets $\bar{J}=1$ for all travelling waves, irrespective of their wave number. For defects, $q_{p}=k_{\infty}$ is a suitable choice. Since defects are sources or sinks of travelling waves in both left and right directions, we expect $\bar{J}$ to vanish in the presence of defects. For small $\alpha$, $\bar{J}$ is close to zero. At large $\alpha$, consistent with travelling states, $\bar{J}\sim 1$. In the vicinity of the transition point $\alpha_{c}$ (grey shaded region), $\bar{J}$ shows large fluctuations as some of the simulations can transition to travelling wave states while others show defects. We verify this by plotting the fraction of simulations $(F_{T})$ that show defect states for $\alpha$ in the shaded region (see Fig. 4(b,inset)). For $\alpha\lesssim 0.4$, the entire ensemble shows defect states whereas for $\alpha>0.6$, all the simulations end up in travelling wave states. From the plots of $\rho_{D}$ and $\bar{J}$, we conclude that the disorder to order transition in 1D occurs at $\alpha_{c}\sim 0.6$, which is in agreement with the crossover threshold $\alpha_{\times}\sim 0.62$ predicted by the wave number selection and the Eckhaus instability. This is in direct contrast to the 2D disorder-to-order transition, which occurs at $\alpha\sim 0.28$ [3], and highlights the role of dimensionality on the phenomenology of the NRCH model.
III.4 Resonances
A careful analysis of the $\rho_{D}$ vs. $\alpha$ plot reveals nontrivial features of the defect configurations. We find that $\rho_{D}$ shows sharp minima at certain values of $\alpha$ which we call “resonances” (see Fig. 4(a)). A visual inspection of the evolution of the defect configurations reveals that at resonances, the defects keep merging even at longer times resulting in smaller defect density. For further analysis, we compute the time it takes for the defect configurations to reach a steady state ($t_{\mathrm{st}}$), where $t_{\mathrm{st}}$ is defined as the time after which no defect merger event takes place and the defect density becomes constant. We find that for the majority of $\alpha$-space, defects attain a steady state relatively quickly, however in the resonance regions, $t_{\mathrm{st}}$ can be order of magnitude higher. Consistent with these observations, we also observe sharp rise in the nearest neighbour separation $d_{\mathrm{min}}$ in the resonance-space. Similar phenomenon, where $d_{\mathrm{min}}$ rises sharply and decreases gradually was also observed in the study of pairwise defect interactions in 2D [4]. It also highlights the rich phenomenology of defect interactions in the NRCH model.
IV Dirichlet and Neumann Boundary Conditions
We now consider the effects of varying boundary conditions on the phenomenology of the NRCH model. As discussed in Section II, in addition to the periodic boundary conditions, we consider Neumann and Dirichlet boundary conditions. Few comments are in order before we discuss the numerical results on N-BC and D-BC. Firstly, travelling waves of the form (3) are incompatible with both N-BC and D-BC, thus we do not expect states with complete polar order to emerge beyond $\alpha_{c}$. Secondly, for both N-BC and D-BC, polar order vanishes at the boundaries, i.e.,
$J(x=0,t)=J(x=L,t)=0~\forall~t,$ (8)
as $\phi$ or $\partial_{x}\phi$ are zero at the boundaries for D-BC and N-BC, respectively. Thus, the boundaries act as additional sources or sinks which make it possible to have any number of sources and sinks in the inner domain. In our analysis, we ignore the defects on the boundaries and only consider defects that are of the form described in Fig. 2 and are located inside the domain within $x\in(x_{\mathrm{pad}},L-x_{\mathrm{pad}})$. For both N-BC and D-BC, we heuristically choose $x_{\mathrm{pad}}=50$, which we find, filters the boundary defects well.
Figure 5: Comparison of $k_{l}$ for different $\alpha$ for $\alpha<\alpha_{c}$ for different boundary conditions. For a given $\alpha$, $k_{\infty}$ is higher for both N-BC and D-BC when compared to the $k_{\infty}$ for the P-BC. As a check, we also plot the $k_{\infty}$ obtained from Dedalus for the P-BC, which matches with our own simulations (dashed black line). $\alpha_{\times}=0.35$.
Figure 6: Phenomenology for the N-BC, which is similar to the P-BC for $\alpha<\alpha_{c}$. (a) Plot of average defect density $\rho_{D}$ versus $\alpha$, which matches very well with the statistics for P-BC. (b) Plot of average polar order. For $\alpha<\alpha_{c}$, we find $\bar{J}\sim 0$. For $\alpha>\alpha_{c}$, the system shows partial order. Near the transition point, we observe large intermittent fluctuations, which are reflected in the large fluctuations in $\bar{J}$ as well. (c) Comparison of different length scales. For $\alpha<\alpha_{c}$, $k_{\infty}$ is the dominant length scale and all other scales converge to it. Near the transition, we find large disagreements as well as large fluctuations and different length scales, which once again converge at larger $\alpha$.
Figure 7: Phenomenology for the D-BC. (a) Plot of average defect density $\rho_{D}$ versus $\alpha$, which is in agreement with the P-BC results for $\alpha>0.3$ For smaller $\alpha$, $\rho_{D}$ is significantly lower, and $\rho_{D}$ often vanishes for these $\alpha$. (b) Plot of average polar order. For $\alpha<\alpha_{c}$, when $\rho_{D}>0$, we once again find $\bar{J}\sim 0$. However, certain realizations for $\alpha\in(0.1,0.3)$ reach defect free configurations and thus show significant polar order (orange markers). For $\alpha>\alpha_{c}$, the phenomenology is similar to the N-BC; system shows partial order and near the transition point, we once again observe large fluctuations in $\bar{J}$. (c) Comparison of different length scales which show similar behaviour as N-BC.
Figure 8: Polar order $J$ for three different configurations with no defects, one defect, and multiple defects for $\alpha=0.2$ for D-BC. Markers show the location of defects inside the domain. Since, $|J|>0$ in the entire domain for defect-less configuration, $\bar{J}$ is large.
Figure 9: Kymographs of $J(x,t)$ for N-BC and D-BC for slightly above the transition point ($\alpha=0.70)$ and at higher values of nonreciprocity ($\alpha=1.0$). Closer to the transition point, we observe large transient and intermittent patches of polar order in the domain that keep emerging and vanishing. For large $\alpha$, both N-BC and D-BC show that the system develops two partitions with polar order in opposite directions. These partitions sustain over time but fluctuate around their mean value.
IV.1 Defects below $\alpha_{c}$
Below $\alpha_{c}$, we once again find defect configurations. As shown in Fig. 5, similar to the P-BC, for both N-BC and D-BC, defects with a particular wave number are selected. However, wave number selection is affected by the presence of boundaries; for a given $\alpha$, the selected wave number is higher than that for the P-BC and a least-square fit yields $k_{\infty}=\alpha^{0.52}$. Thus, one expects that defect solutions will vanish around $\alpha\sim 1$. It is tempting to use the Eckhaus instability to arrive at a crossover threshold of nonreciprocity; $\alpha_{\times}\sim 0.33$ for N-BC and D-BC. However, we note that the instability analysis assumes travelling wave states, which are not permitted solutions for these boundary conditions. Our numerical analysis suggests that the transition occurs at $\alpha_{c}\sim 0.6$.
The phenomenology of the NRCH model with N-BC is qualitatively similar to the P-BC. For N-BC, the defect density $\rho_{D}$ agrees with the results from the P-BC (see Fig. 8(a)). Further, the average polar order vanishes for defect states. However, D-BC show a distinct phenomenology as compared to the two. For $\alpha\in(0.1,0.2)$, the system quickly attains a steady-state configuration, which can have any number of defects, including no defects at all. Thus the defect density shows large fluctuations for these values (see Fig. 8(a)). Since, the system can end up in defect-free or defect-laden states for D-BC (see Fig. 8), while computing the average polar order, we make a distinction between the two kind of states and find that as expected defect-laden configuration show negligible polar order. On the other hand, defect-free configurations exhibit large average polar order. For these states, one boundary acts as a source, and the other acts as a sink. Irrespective of the number of the defects remaining at the end, the system exhibits wave number selection and the selected wave number is the same. This is expected as the defect free state is achieved by consecutive merger of defects, and thus the wave number is selected prior to the formation of this state. For $\alpha\geq 0.3$, $\rho_{D}$ agrees with the results for P-BC.
IV.2 Fluctuating polar order beyond $\alpha_{c}$
Since travelling waves are only permitted for P-BC, we expect the phenomenology for $\alpha>\alpha_{c}$ to vary significantly for N-BC and D-BC. Indeed, beyond $\alpha_{c}$, the system does not attain a single travelling wave phase. For $\alpha\gtrsim\alpha_{c}$, both N-BC and D-BC show intermittent behaviour. The emergence of global polar order is thwarted by the incompatibility of travelling bands with the boundary conditions. Thus, the system does not settle down, but shows transient, fluctuating states with patchy polar order that sustain over long times (see Fig. 9).
In the case of P-BC, a single wavenumber dominates the steady states, for defect configurations below $\alpha_{c}$ it is the selected wavenumber $k_{\infty}$, for the travelling states above $\alpha_{c}$, it is the wave number of the travelling wave. For N-BC and D-BC, we find that the intermittent patchy ordered states exhibit fluctuations at multiple wave numbers. To verify this, we compute various relevant scales using the structure factor $S(k,t)$: (i) The dominant mode $k_{\mathrm{d}}$ defined as the wavenumber at which $S(k)$ peaks, i.e, $k_{\mathrm{d}}=\mathrm{argmax}_{k}S(k)$, (ii) the coarsening mode $k_{\mathrm{c}}=\sum_{k}kS(k)/\sum_{k}S(k)$, and (iii) the integral mode $k_{\mathrm{i}}=\sum_{k}S(k)/\left(\sum_{k}S(k)/k\right)$. Here, $S(k,t)=|\psi(k)|^{2}$ is the structure factor, where $\psi(k,t)=\sum_{x}\psi(x,t)\exp(-ikx)$ is the Fourier transform of the windowed field $\psi(x,t)=w(x)\phi(x,t)$. Note that we apply the Hanning window $w(x)=\mathrm{sin}^{2}(\pi x/L)$ on the data for the N-BC and D-BC to avoid spurious amplitude errors arising from the aperiodicity. As shown in Fig. 8(c) and Fig. 8(c), $k_{\mathrm{d}}$, $k_{\mathrm{c}}$, and $k_{\mathrm{i}}$ are in excellent agreement with each other and match with $k_{\infty}$ for $\alpha<\alpha_{c}$, emphasizing that $k_{\infty}$ governs the dynamics for defect configurations. Slightly above $\alpha_{c}$, different modes disagree with each other, and exhibit large fluctuations, signifying that the transient states show fluctuations at multiple length scales. Consequently, it is not possible to choose a suitable normalization for the average polar order; i.e., a suitable choice for $q_{p}$ in Eq. 7 is not available for these states, which is also reflected in large fluctuations of $\bar{J}$ (see Fig. 8(b) and Fig. 8(b)) that were computed using $k_{\mathrm{c}}$.
For $\alpha>\alpha_{c}$, the system solves the incompatibility with the travelling waves by creating two domains with opposite polar order (see Fig. 9). The system boundary and the partition between the two domains with opposite polar order acts as sources or sinks of the polar order. The partition can form anywhere inside the domain and fluctuates slowly with time. In this case, the distribution function of $J(x,t)$ is primarily bimodal, where the strength of the two peaks depends the location of partition in the domain. As a consequence, we find large non zero polar order when the partition is closer to one of the domains, or small polar order if it close to the middle of the simulation domain. Finally, various scales, $k_{\mathrm{d}}$, $k_{\mathrm{c}}$, and $k_{\mathrm{i}}$, all converge to similar values.
V Conclusions
In this paper, we have presented a systematic study of the disorder to order transition for the 1D NRCH model. In 1D, defect solutions are the sources and sinks of travelling waves. Sources are analogous to the two-dimensional topologically neutral targets, whereas sinks are similar to the disinclination lines where the waves emanated from sources meet. For a given $\alpha$, defects select a unique wave number $k_{\infty}$ that increases monotonically with $\alpha$. Wave number selection predicts a crossover threshold $\alpha_{\times}\sim 0.62$ above which the defect solutions cannot exist.
At small $\alpha$, disordered initial states evolve into multi-defect configurations, where sources and sinks are arranged in an alternating manner. With increasing $\alpha$, overall the defect density increases, peaks at $\alpha\sim 0.2$, and then vanishes at $\alpha_{c}\sim 0.6$, which marks the onset of disorder to order transition. In accordance with this observation, we find that there is no significant average polar order for $\alpha<\alpha_{c}$, but travelling states for $\alpha>\alpha_{c}$ show perfect order in the long time steady state. In a direct contrast to 2D, where the transition occurs at $\alpha_{c}\sim 0.28\ll\alpha_{\times}\sim 0.58$, the transition point in 1D, $\alpha_{c}\sim 0.6$ is in close agreement with the crossover threshold $\alpha_{\times}\sim 0.62$ predicted by wavenumber selection. A closer inspection of the defect dynamics reveals that at certain resonance values of nonreciprocity below $\alpha_{c}$, $\rho_{D}$ shows sharp minima, which is a consequence of the fact that for these values of $\alpha$, defect merger events continue to occur even at very long times. Additionally, the average inter-defect separation rises sharply for these $\alpha_{c}$, which was also observed in the study of pairwise defect interactions in 2D [4].
Our numerical simulations with N-BC and D-BC, which are inconsistent with the travelling waves, show different flavours of the disorder-order transition. While the phenomenology for the N-BC is similar to the P-BC for $\alpha<\alpha_{c}$, D-BC significantly affects the defect dynamics and below $\alpha_{c}$, we find configurations with one, many, or no defects at all. Above $\alpha_{c}$, N-BC and D-BC share similar phenomenology, which is distinct from the P-BC. For nonreciprocity slightly above the threshold, we find intermittent, fluctuating polar order. For larger values of $\alpha$, the system partitions into two subdomains which sustain over time.
To conclude, our study shows that for the P-BC NRCH model shows similar phenomenological behaviour in both 1D and 2D, with a crucial difference that the disorder-to-order transition occurs at the crossover value predicted by wave number selection in 1D. We further show that boundary conditions can significantly affect the phenomenology. These findings call for future studies to further investigate the defect dynamics and defect interactions for conserved systems with nonreciprocal interactions.